OpenStudy (anonymous):

solve this equation for x 4^x =30

OpenStudy (anonymous):

\[x=\log_{4} 30\] after I cant remember how to deal with

OpenStudy (anonymous):

after that

OpenStudy (anonymous):

\[4^x = 30 \implies ln 4^x = ln 30 \implies x ln 4 = ln 30 \implies x = \frac{ln 30}{ln 4}\]

OpenStudy (anonymous):

ttt is also correct. \(\frac{ln 30}{ln 4} = log_4 30\)